Question 5

You are considering a project with the following after-tax cash flows (in millions of dollars):

Years from Now           After-tax CF

            0                      -50      

            1-4                   10

            5                      30   

Suppose the project is only financed by equity and its beta is 1.2. Given Rf = 3% and E(RM) = 10%, you’d use the CAPM to estimate the (equity) cost of capital.

What is the NPV of the project? ______ millions
Answer #1

-SUM(G3:G8) D Е F G Н 1 pv@11.4% | Present value 1.0000 $ 2 year Cash flows (50.00) 0S (50.00) 1$ 0.8977 $ 4 10.00 8.98 0.805

f SUM(G3:G8) G9 Е F G Н 1 Cash flows pv@11.4% Present value 2 year 3 о |-ЕЗ*ЕЗ -50 1 - F3/ 1.114 =F4/1.114 |=F5/1.114 =F6/1.1